Help me with the workings to these questions please!!!

Question 4 d when a 8.9, b 13.0 and c 7.8 will equal (correct to four decimal places) Not answered Marked out of 1.00 Flag question 4cr Answer: The correct answer is:0.0464 Question 5 Not answered Marked out of 1.00 ch For 3ry 2v-4,find at x 4.0 andy 4.3 (correct to four decimal places) Flag question Answer: The correct answer is: -1.6145 Question 6 Not answered Marked out of 1.00 units are produced and b 66.24 is (correct to two decimal places) Demand for a product is given by the function p -b-48 where p ls price and q is quantty. The value of Marginal Revenue when 66 Flag question Answer: The correct answer is: -1878.84
Answer #1

1. large int_{a}^{b}rac{1}{4+cx}dx

Where a= 8.9, b = 13 and c= 7.8

Let us assume 4+7.8x = u

Now differentiate u wrt x we get

large rac{du}{dx}= rac{d(4+7.8x)}{dx}

large rac{du}{dx}= 7.8

large rac{du}{7.8}= dx

When a= 8.9 then u = 4 + 7.8*8.9 = 73.42

When b = 13, u = 4+ 7.8*13 = 105.4

Now plug in this into the main equation.

large int_{a}^{b}rac{1}{4+cx}dx

= large int_{73.42}^{105.4}rac{1}{7.8u}du

=large rac{1}{7.8}*int_{73.42}^{105.4}rac{1}{u}du

=large rac{1}{7.8}*lnleft | u ight |_{73.42}^{105.4}

= (1/7.8)*[ln(105.4) -ln(73.42)]

=(1/7.8) *(4.657762 - 4.296196)

=(1/7.8)*0.361566

= 0.046354

=0.0464 (Approximately)

2. Given, large 3x^3y^2 = 2y-4

Differentiating above equation wrt x we get

dr dr

large 9x^2y^2 + 6x^3yrac{dy}{dx}=2 rac{dy}{dx}

large 9x^2y^2 =(2- 6x^3y)rac{dy}{dx}

dy9x2y 2

Plug in the value of x and y

dy9* 424.32 dz 2-6*43 * 4.3

dy2662.56 da 2 -1651.2

2662.56 1649.2 dy

dy^ -1.614455

Approximately the answer is -1.6145

3. Given large P = b - 48q^{rac{3}{4}}

Total Revenue = PQ

large TR = Pq =( b - 48q^{rac{3}{4}})q

TR = bg-48g(14)

large TR = bq - 48q^{rac{7}{4}}

Differentiating above equation wrt q we get

large MR = b - 48*rac{7}{4}*q^{rac{3}{4}}

Now plug in b = 66.24, q = 66

MR 66.24-48 *-* 667

MR = - 1878.84

Please contact if having any query thank you. Kindly rate up will be obliged to you for your generous support.